OpenStudy (anonymous):

The formula for voltage (in volts) is given by V = IR, where I is the current in amperes, and R is the resistance in ohms. If a light bulb uses 0.4 amperes of current at a voltage of 1.8 volts, the formula to find R is , and the resistance of the bulb is ohms.

OpenStudy (astrophysics):

\[V = I R\] we're just solving for R here, so manipulate the variables here and solve for R, any idea?

OpenStudy (astrophysics):

See I is being multiplied by R so all you have to do is use the opposite operation of what already exists to both sides, which would be...?

OpenStudy (anonymous):

.4r=1.8 right?

OpenStudy (astrophysics):

Yes, now solve for r

OpenStudy (anonymous):

so r=4.5

OpenStudy (astrophysics):

Yup, \[R = \frac{ V }{ I }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

R=V/I=1.8V/0.4A=9/2ohm=4,5ohm You have Ohm's Law : ( I in amps ) = ( V in volts ) / ( R in ohms ) R in ohms = ( V in volts ) / ( I in amps ) R in ohms = ( 1.8 volts ) / ( 0.4 amps ) R in ohms = 4.5 ohms

OpenStudy (denisaboichuk):

32